A thread charged homogeneously with line density λ=3.47 μC/m forms a loop in the shape of a cut circle with radius r=6.12m, cut by a line parallel to the y-axis with length l=10.54. Calculate the elctric field at the center of the circle. State direction and magnitude of the electric field vector.
1
Expert's answer
2019-04-22T11:09:03-0400

Charge line density "\lambda = 3.47 \, \mu\text{C/m}"

Radius of circle "r = 6.12 \, \text{m}"

Length of cut "L = 10.54 \, \text{m}"

Lets divide arc by small parts with angle "d\varphi" . It has length "dl = rd\varphi" and charge "dq = \lambda r d\varphi" , therefore produces electric field with magnitude


"dE_y = - \frac{1}{4\pi\varepsilon_0}\frac{\lambda r d \varphi}{r^2} \sin \varphi, \, dE_y = - \frac{1}{4\pi\varepsilon_0}\frac{\lambda r d \varphi}{r^2} \cos \varphi"

As electromagnetic field obeys superposition principle, therefore we integrate in ranges from "-\theta" to "\theta" :



"E_x = - \frac{1}{4\pi\varepsilon_0}\int_{-\theta}^{\theta} \frac{\lambda r d \varphi}{r^2} \sin \varphi = \frac{\lambda}{4\pi\varepsilon_0 r} \cos \varphi |_{-\theta}^{\theta} = 0"

"E_y = - \frac{1}{4\pi\varepsilon_0}\int_{-\theta}^{\theta} \frac{\lambda r d \varphi}{r^2} \cos \varphi =- \frac{\lambda}{4\pi\varepsilon_0 r} \sin \varphi |_{-\theta}^{\theta} = - \frac{\lambda}{2\pi\varepsilon_0 r} \sin \theta"

where θ defines as

"sin \theta = \frac{L}{2r}"

Therefore, magnitude of electric field equals to abs. value of y component and directed parallel to y-line opposite to arc.


"E_y = - \frac{\lambda L}{4\pi\varepsilon_0 r^2},\, E_x=0"

"E=\frac{\lambda L}{4\pi\varepsilon_0 r^2}= 8788 \text{V/m}"